# Explorations of Self-Repair in Language Models

### v1

@@ 5.1 Implications of Imperfect Self-Repair for Interpretability Efforts
In practice, this is most concerning when considering circuit analysis, where we patch or ablate individual model components and observe their effect on the output, in order to isolate the sparse subgraph of the model relevant to our task, as performed in (Wang et al., 2023; Lieberum et al., 2023; Conmy et al., 2023). However, fortunately, circuit analysis only requires identifying whether a component is important or unimportant (i.e. whether it belongs in the sparse subgraph), not the precise effect of ablating it. So the fact that self-repair is imperfect across the general distribution (Section 2.2) helps reduce some of the concerns for circuit discovery efforts; because the ’importance’ of various components is extremely heavy-tailed, even a significant fractional decrease in the estimated effect won’t change which nodes are important.
@@
5.2 Taking Models Off-Distribution can Cause Unexpected Consequences that are Easy to Misinterpret
@@
When we first found self-repair, we assumed that models were intentionally self-repairing. However, the LayerNorm-based self-repair seems to be a side effect of unrelated mathematical properties of the model, interacting with how our ablations change the norm of the residual stream. It’s possible that many observed instances of ’self-repair’ are just uninteresting consequences of throwing the model off-distribution, for which this will be clear after discovering a few more insights like the above.
@@ 5.3 Iterative Inference Hypothesis
One perspective from which self-repair becomes less surprising is what we call the Iterative Inference hypothesis: rather than layers being part of a top-down process, assisting all future layers in complex circuits, many language model components are more of a bottom-up process, where each layer treats the input as a guess for the final output, and tries to reduce the error between this guess and the true next token.
@@
This is consistent with the evidence presented in McDougall et al. 2023 on how specific heads can influence downstream heads to not perform a task. Another line of evidence that supports this hypothesis is the identification of attention heads across models which we dub ’self-reinforcing’ and ’self-repressing’.
@@
We observe ’self-reinforcing’ and ’self-repressing’ heads, in which the direct effect of the head increases or decreases proportional to how much of the original output of the head is added. Figure 9 highlights one head in GPT2-Medium which ’self-represses’. The repression is proportional to the amount the output is scaled (Appendix E).
@@ 5.3.1 Additional Discussion on the Iterative Inference Hypothesis
Additionally, when moving the residual stream read by the "Name Movers" into the residual stream read by the "Backup Name Movers’, the backup heads begin performing name moving as well (note that this is equivalent to zero ablating all the attention heads and MLP layers in between them). This further highlights the likelihood of there being a shared stimulus to perform Name Moving which both types of heads respond to, but which the Name Movers often respond to and ablate.
However, there is also a wealth of evidence against this hypothesis. Anecdotally, previous token and induction heads (Olsson et al., 2022) don’t seem to be self-repaired. Perhaps this may be because these heads are always expected to perform these ’fundamental’ tasks across all prompts. Further, prior literature has emphasized the presence of complex circuits within the model, which don’t easily exist across layers, the S-inhibition heads in the Indirect Object Identification task (Wang et al., 2023) being one of them.
This highlights how specific tasks in the head may not be performed ’iteratively’: they may be too fundamental such that the model has specific heads dedicated to them, or it may be too complex/costly to perform iteratively by multiple heads. Many heads in the model may not be performing ’iterative’ tasks.
@@ 6 Related Work
The self-repair phenomena was first identified in the Indirect Object Identification distribution (Wang et al., 2023). Self-Repair was initially explored as the Hydra Effect (McGrath et al., 2023), and a specific instance of it was explained by Copy Suppression (McDougall et al., 2023).
@@ 7 Conclusion
A strong limitation of our work is that it focuses on the repairing of direct effects as opposed to other ’behavioral’ measures of attention heads (which aren’t immediately represented as direct effects). It’s unclear how much the mechanisms highlighted in sections 3 and 4 will generalize to these other potentially viable forms of self-repair. Despite this, we expect that the ideas presented here will help provide some grounding for further research on self-repair.
@@ Appendix B Mathematical Derivation of LayerNorm Self-Repair
How can we explain this logit difference? As argued in Elhage et al. 2021, we can simplify the LayerNorm operation such that we ’fold in’ LayerNorm projections to the weights of linear layers before and after the projection, leaving the nonlinear operation of scaling the residual stream by dividing by a scaling factor proportional to the norm of the residual stream.
@@ Appendix D Residual Stream Norms Change when Ablating
Appendix E Self-Repressing Heads
The existence of Self-Repressing heads (Section 5.3) helps support the Model Iterativity hypothesis because they robustly respond to their own output. However, one interesting aspect of this is that the amount in which these heads self-repress is proportional to the amount in which scale the heads output back into itself. Figure 19 highlights the self-repressing head L21H1 in GPT2-Medium, but plots the forward passes when adding in the output of the head scaled by 1, 3, and 5 times the existing output.
@@ Appendix E Self-Repressing Heads
Appendix F Self-Repair Graphs Across Different Models.
In Figures 21 through 22, we plot the self-repair of individual attention heads for different models, similar to the experimental setup for Figure 2.
Figure 20: Llama-7b self-repair, per head

### v2

@@ 5.1 Implications of Imperfect Self-Repair for Interpretability Efforts
In practice, this is most concerning when considering circuit analysis, where we patch or ablate individual model components and observe their effect on the output, in order to isolate the sparse subgraph of the model relevant to our task, as performed in (Wang et al., 2023; Lieberum et al., 2023; Conmy et al., 2023). However, fortunately, circuit analysis only requires identifying whether a component is important or unimportant (i.e. whether it belongs in the sparse subgraph), not the precise effect of ablating it. So the fact that self-repair is imperfect across the general distribution (Section 2.2) helps reduce some of the concerns for circuit discovery efforts; because the ‘importance’ of various components is extremely heavy-tailed, even a significant fractional decrease in the estimated effect won’t change which nodes are important.
@@
5.2 Taking Models Off-Distribution can Cause Unexpected, Easily Misinterpretable, Consequences
@@
When we first found self-repair, we assumed that models were intentionally self-repairing. However, the LayerNorm-based self-repair seems to be a side effect of unrelated mathematical properties of the model, interacting with how our ablations change the norm of the residual stream. It’s possible that many observed instances of ‘self-repair’ are just uninteresting consequences of throwing the model off-distribution, for which this will be clear after discovering a few more insights like the above.
@@ 5.3 Iterative Inference Hypothesis
One perspective from which self-repair becomes less surprising is what we call the Iterative Inference hypothesis: rather than layers being part of a top-down process, assisting all future layers in complex circuits, many model components are more of a bottom-up process, where each layer treats the input as a guess for the final output, and tries to reduce the error between this guess and the true next token.
@@
5.3.1 Evidence for, and Discussion of the Iterative Inference Hypothesis
What supports the Iterative Inference Hypothesis? Firstly, the hypothesis is consistent with the evidence presented in McDougall et al. 2023 on how specific heads can influence downstream heads to not perform a task. Another line of evidence that supports this hypothesis is the identification of attention heads across models which we dub ‘self-reinforcing’ and ‘self-repressing’.
@@ 5.3.1 Evidence for, and Discussion of the Iterative Inference Hypothesis
We observe ‘self-reinforcing’ and ‘self-repressing’ heads, in which the direct effect of the head increases or decreases proportional to how much of the original output of the head is added. Figure 9 highlights one head in GPT2-Medium which ‘self-represses’. The repression is proportional to the amount the output is scaled (Appendix F).
@@
Additionally, when moving the residual stream read by the "Name Movers" into the residual stream read by the "Backup Name Movers", the backup heads begin performing name moving as well (note that this is equivalent to zero ablating all the attention heads and MLP layers in between them). This further highlights the likelihood of there being a shared stimulus to perform Name Moving which both types of heads respond to, but which the Name Movers often respond to and ablate.
However, there is also a wealth of evidence against this hypothesis. Anecdotally, previous token and induction heads (Olsson et al., 2022) don’t seem to be self-repaired. Perhaps this may be because these heads are always expected to perform these ‘fundamental’ tasks across all prompts. Further, prior literature has emphasized the presence of complex circuits within the model, which don’t easily exist across layers, the S-inhibition heads in the Indirect Object Identification task (Wang et al., 2023) being one of them.
This highlights how specific tasks in the head may not be performed ‘iteratively’: they may be too fundamental such that the model has specific heads dedicated to them, or it may be too complex/costly to perform iteratively by multiple heads. Many heads in the model may not be performing ‘iterative’ tasks.
@@ 6 Related Work
The self-repair phenomena was first identified in the Indirect Object Identification distribution (Wang et al., 2023). Self-Repair was initially explored as the Hydra Effect (McGrath et al., 2023), and a specific instance of it was explained by Copy Suppression (McDougall et al., 2023). Our work differs from this prior work as follows:
  1. 1.
    We analyze self-repair across the a pretraining distribution and with individual attention heads. Past work has mostly highlighted the presence of self-repair in narrow distributions or when ablating larger components.
  2. 2.
    Our introduction of LayerNorm is as a self-repairing mechanism is an entirely novel contribution.
  3. 3.
    Prior work has identified self-repairing Anti-Erasure in entire MLP layers (McGrath et al., 2023) or discussed Anti-Erasure in the context of Suppression heads (McDougall et al., 2023). Our work highlights how erasure in MLP layers is better understood in terms of ‘erasure neurons’, which are sparse and self-repair significantly.
@@
A related phenomena is parameter redundancy, focused on reducing parameters without harming performance. Self-repair is a different perspective, but may explain how one can ablate individual attention heads without significant performance loss, as demonstrated in (Michel et al., 2019).
@@ 7 Conclusion
A strong limitation of our work is that it focuses on the repairing of direct effects as opposed to other ‘behavioral’ measures of attention heads (which aren’t immediately represented as direct effects). It’s unclear how much the mechanisms highlighted in sections 3 and 4 will generalize to these other potentially viable forms of self-repair. Despite this, we expect that the ideas presented here will help provide some grounding for further research on self-repair.
Impact Statement
This paper aims to advance the field of Artificial Intelligence interpretability. We hope these insights help build tools to make uninterpretable machine learning systems more interpretable, and thus more understandable, safe, and controllable.
@@ Appendix B Mathematical Derivation of LayerNorm Self-Repair
How can we explain this logit difference? As argued in Elhage et al. 2021, we can simplify the LayerNorm operation such that we ‘fold in’ LayerNorm projections to the weights of linear layers before and after the projection, leaving the nonlinear operation of scaling the residual stream by dividing by a scaling factor proportional to the norm of the residual stream.
@@ Appendix D Residual Stream Norms Change when Ablating
Appendix E Centered Logits
An important aspect of our analysis is that the logits at each position are centered such that the average logit is zero. This helps control for changes in logits which can be explained by translations in all the logits - which is important, because a translation in logits leads to no change in the final log probabilities of the outputs.
As adding a constant to the logits does not affect the output probabilities (softmax is translation invariant), centered logits are more amenable to analysis. As such, we cannot measure parts of self-repair coming from adding a constant to the logits of all tokens simultaneously, as the centering of the logits would remove this behavior. Instead, the observation of self-repair in a logit - such as in Figure 5 - indeed reflects a self-repair in the correct token (and not all tokens).
Appendix F Self-Repressing Heads
The existence of Self-Repressing heads (Section 5.3) helps support the Iterative Inference hypothesis because they robustly respond to their own output. However, one interesting aspect of this is that the amount in which these heads self-repress is proportional to the amount in which scale the heads output back into itself. Figure 19 highlights the self-repressing head L21H1 in GPT2-Medium, but plots the forward passes when adding in the output of the head scaled by 1, 3, and 5 times the existing output.
@@ Appendix F Self-Repressing Heads
Appendix G LayerNorm Ratios are Significant
For many models, the direct effect of the head is often ‘significant’ when its direct effect is greater than 1 (it can get to up to 4, but on rare occasions). However, the difference between two logits can be up to around 10 or more logits! For a ratio such as 1.02, a difference of 10 logits between two tokens would lead to a difference of 0.2 logits in how much LayerNorm self-repairs one token over another in the toy example presented in Section 3 (the amount a token A is self-repaired relative to another token B can be calculated via $(\frac{S}{S\textquoteright}-1)(logit_{a}-logit_{b})$, where $logit_{a}$ and $logit_{b}$ are the final logits of A and B - this equation can be easily derived from the first presented equation in Section 3). This is a nontrivial fraction of a direct effect.
Appendix H Self-Repair Graphs Across Different Models.
In Figures 21 through 22, we plot the self-repair of individual attention heads for different models, similar to the experimental setup for Figure 2. We also highlight what idealized versions of ‘perfect self-repair’ and no self-repair my look like in Figures 24 and 24.
@@ Appendix G LayerNorm Ratios are Significant
Figure 23: An idealized version of perfect self-repair
Figure 24: An idealized version of no self-repair